# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v17

@@ 10.4 Trigonometric Library Errors
Figure 24: Difference between library and Quart $\sin(x)$ (y-axis) for $x=2\pi j/2^{L},j=0,1\dots 2^{L+2}$ (x-axis), and $L=5,6$ (legend). The uncertainties of the Quart $\sin(x)$ is $\sin(x)$ ULP, which shows a periodicity of $\pi$.
@@
Formula (10.1) and (10.2) restrict the use of $\sin(x)$ and $\cos(x)$ to $x=2\pi j/2^{L}$, where $L$ is the FFT order. To minimize numerical errors in computing $\sin(x)$, the following indexed sine can be used in place of standard library sine functions:
@@
    The values of $\sin(\pi j/2^{L}),j\in[0,2^{L-2}]$ are library sine directly, while the values of $\sin(\pi j/2^{L}),j\in[2^{L-2},2^{L-1}]$ are computed from library $\cos(\pi(2^{L-1}-j)/2^{L})$.
@@
Because the Quart sine function strictly preserves the symmetry and periodicity of sine function, it provides superior numerical accuracy compared to the Library sine function.
@@
  • Figure 25 shows the value difference between the Quart and Library $\cos(x)/\sin(x)$ also increases roughly linearly with $|x|$ but are $10^{2}$ times larger than those observed for $\sin(x)$. Therefore, the linear spectrum in Formula (10.3) contains significantly larger numerical errors when computed using library sine functions.
@@ 10.5 Using Quart Sine for Sin/Cos Signals
Figure 26: Histograms of normalized errors of Sin/Cos signals for forward, reverse and roundtrip transformations (legend) using the Quart sine function. The FFT order is $18$.
With FFT order as the specific dimension, the error deviations for forward and reverse transformations resemble Figure 14, while those for round-trip transformation is almost identical to Figure 15, independent of Sin or Cos signals, or frequency of the signals. Therefore, the results for Sin and Cos signals across all frequencies are pooled together for statistical analysis, under the unified category Sin/Cos signals. When the FFT order $L$ is less than 8, the error deviations oscillate around $1$ due to insufficient sample count of $2^{L}$. Even data for forward and reverse transformations are drastically different, variance arithmetic works effectively in all cases.
When $L=18$ and $\delta x=0$, Figure 26 shows that the error distributions of Sin/Cos signal resemble Normal distribution, with an additional Delta-like distribution at $\tilde{z}=0$. The error distribution of the reverse transformation is structured on top of the Normal distribution, suggesting that the reverse transformation is more sensitive to numerical errors in sine function.
@@ 10.6 Using Library Sine for Sin/Cos Signals
Figure 27: Histograms of normalized errors of Sin/Cos signal for forward, reverse and roundtrip transformations (legend) computed using the Library sine function. The FFT order is $18$.
@@
Figure 28: FFT value error spectrum of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the forward transformation. The legend distinguishes between uncertainty and value error. The x-axis represents the frequency index, while the y-axis represents both uncertainty and value error.
@@
Figure 29: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the reverse transformation. The legend distinguishes between uncertainty and value error. The x-axis represents the time index, while the y-axis represents both uncertainty and value error.
@@
Figure 30: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the roundtrip transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the frequency index, while the y-axis represents both uncertainty and value error.
@@
Figure 31: Error deviation (z-axis) of FFT reverse transformation of $sin(f\frac{2\pi}{2^{L}}j)$ versus frequency $f$ (x-axis) and FFT Order $L$ (y-axis).
@@
With FFT order as the specific dimension, the error deviations for the forward and the reverse transformations resemble Figure 14, while those for the round-trip transformation is almost identical to Figure 15. In addition, the error deviations are larger than $1$ when $\delta x<10^{-15}$ for the forward transformation and $\delta x<10^{-14}$ for the reverse transformation.
@@
When $\delta x=0$, the error deviations for the reverse transformation increase with FFT order, to $6.2$ at FFT order $18$. As shown in Figure 24, the Library sine function contains more numerical errors, so the error distribution for the reverse transformation using Library sine function in Figure 27 is more structured and broader than that in Figure 26 using Quart sine function, while the error distribution for the forward transformations is more similar. This difference is consistent with a larger error deviation $6.2>1$ for the reverse transformation, but a comparable error deviation $1.1>1$ for the forward transformation when $\delta x=0$.
@@
Using the Library sine function, for a sine wave with a frequency of $3$, Figure 28, 29, and 30 presents the value errors for the forward, the reverse, and the round-trip transformations, respectively. In the reverse transformation, value errors exhibit a clear trend of increasing with the time index. These large value errors appear systematic rather than random and visually resemble a resonant pattern. Similar increases are observed at other frequencies and FFT orders, as well as in computational results obtained using mathematical libraries such as SciPy. In contrast, such resonance is absent for the roundtrip transformation as shown in Figure 30, as well as when using the Quart sine function. Figure 31 demonstrates that the error deviations increase with sine or cosine frequency. Figure 24 indicates that the numerical errors using the Library sine function increase with a periodicity of $\pi$, which may resonate with a signal whose periodicity is an integer multiply of $\pi$, producing the resonant pattern in Figure 29. At higher frequency, the resonant beats between the signal and the numerical errors in the Library sine function become stronger. To suppress these numerical error resonances, an input noise of $\delta x=10^{-14}$ must be added to the sine or cosine signals. Such resonance of numerical errors can easily and mistakenly be taken as signals.
10.7 Using Quart Sine for Linear Signal
@@
Figure 34: Histograms of normalized errors of Linear signals for forward, reverse and roundtrip transformations (legend) computed using the Quart sine function. The FFT order is $18$.
Figure 32 and 33 show the error deviation for the forward and the reverse transformations, respectively. The forward transformation exhibits a larger ideal coverage area than the reverse transformations: $\delta x>10^{-12}$ for the forward transformation, and $\delta x>10^{-8}$ for the reverse transformation. In other areas, both transformations achieve proper coverage with error values around $1$.
When $L=18$ and $\delta x=0$, the error distribution of the reverse transformation in Figure 34 is narrower than that in Figure 27. The corresponding error deviations are $1.5<6.2$, respectively.
10.8 Using Library Sine for Linear Signal
Figure 35: Histograms of normalized errors of Linear signals for forward, reverse and roundtrip transformations (legend) computed using the Library sine function. The FFT order is $18$.
@@ 10.8 Using Library Sine for Linear Signal
Figure 38: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the roundtrip transformations of Linear signals computed using the Library sine function.
Figure 35 shows that the reverse error distribution when $\delta x=0$ is no longer bound. The difference between Figure 35 and Figure 34 is consistent with the large numerical errors as demonstrated in Figure 25. Variance arithmetic fails because of the large amount of unspecified numerical errors from the Library sine.
Figure 36 and 37 shows a much smaller ideal coverage areas than those in Figure 32 and 33. Because result deviations grow more slowly in the reverse transformation than in the forward transformation, the reverse transformation exhibits a smaller ideal coverage region. Outside the ideal coverage region, proper coverage cannot be achieved for the reverse transformation. Furthermore, the range of input noise that produces ideal coverage decreases with increasing FFT order. At sufficiently high FFT orders, visually beyond FFT order $25$ for the reverse transformation, ideal coverage may no longer be achievable. Although FFT is widely regarded as one of the most robust numerical algorithms [5][10], and generally insensitive to input errors, it can still fail due to numerical errors in the Library sine function. Such deterioration in calculation accuracy is not easily detectable when using conventional floating-point arithmetic.
Figure 38 shows that, even when variance arithmetic can no longer effectively track the value errors for the forward and the reverse transformations, it can still effectively track the value errors for the round-trip transformation, as the plateau region at high $L$ and low $\delta x$ in Figure 38. Such error cancellation is due to dependence tracing in statistical Taylor expansion.
@@ 10.9 Ideal Coverage
Figure 39: Result error deviation (left y-axis) and uncertainty deviation (right y-axis) of Linear signals versus FFT order (x-axis) and transformation types (legend) computed using the Library sine function.
Adding noise to the input can suppress unspecified input errors. After adding a Gaussian input noise of $\delta x=10^{-3}$ to a Linear signal when using the Library sine function, the error distributions for both the forward and the reverse transformations become Normal, while the error distribution for the round-trip transformation becomes Delta. Figure 39 illustrates the corresponding error deviations and uncertainty deviations versus FFT order:
@@
  • As expected, the resulting uncertainty deviations for the round-trip transformations remains equal to the corresponding input uncertainties of $10^{-3}$.
@@
  • As expected, the resulting error deviations for the round-trip transformations are far less than $1$ but increase with FFT order $L$ exponentially due to increasing calculation.
@@
Table 2 shows the minimal required noise to achieve ideal coverage for FFT transformations at FFT order $L=18$ for different signals and sine functions, which is consistent with the corresponding error distributions in Figure 26, 27, 34, and 35. This shows that an error distribution can qualify its input uncertainty quantification. Without knowing precise result, similar histogram can be constructed from the calculated mean $\overline{f}$ and deviation $\delta f$, and the result data set for $f$. It is worth investigating if such empirical histogram has similar power to reveal tracking of value errors by variance arithmetic.
@@ 11.1 Summary
When uncorrelated uncertainty condition is met, statistical Taylor expansion outputs the mean, deviation, and reliability of an analytic expression. It tracks the variable dependencies in intermediate steps, rejects invalid calculations, and explicitly incorporates the sample counts and the input uncertainty distributions into the statistical analysis. While statistical Taylor expansion eliminates the dependency problem, it also removes most execution flexibility.
The presence of ideal coverage is a necessary condition for a numerical algorithm using statistical Taylor expansion to be considered correct. The ideal coverage defines the optimal applicable range for an algorithm.
Variance arithmetic simplifies statistical Taylor expansion by introducing numerical rules to elimination invalid results, such as divergent, negative variance, unstable, infinite, or unreliable results. It provides proper coverage for floating-point rounding errors. The applicability of variance arithmetic has been demonstrated across a wide range of computational scenarios.
Library mathematical functions should be recalculated using variance arithmetic, so that each output value is accompanied by its corresponding uncertainty. Without this refinement, the value errors in the library functions can produce unpredictable and potentially significant effects on numerical results.
The code and analysis framework for variance arithmetic are available as an open-source project at https://github.com/Chengpu0707/VarianceArithmetic. A more detailed description of this paper is presented at https://arxiv.org/abs/2410.01223.
@@ 11.2 Improvements Needed
This paper presents statistical Taylor expansion and variance arithmetic, which is still in early stage of development, with several important questions remaining.
@@
Bound momentum $\zeta(2n)$ needs to be extended to all probability distributions. The statistical bounding to find the bounding range $\kappa$ for a sample count needs to be extended to discrete distributions. The choice of ideal bounding range $\hat{\kappa}$ needs to extend to other distributions. The statistical meaning for ideal ratio $\alpha$ in the context of its choice of ideal bounding range $\hat{\kappa}$ needs further investigation: Is $\alpha$ the possibility for the result to be bound by $\hat{\kappa}$?
@@
The performance of variance arithmetic must be improved for broader practical adoption. The fundamental formulas of statistical Taylor expansion, Formula (2.5), (2.3), (2.8), and (2.3), contain a large number of independent summations, making them excellent candidates for parallel processing. Moreover, the inherently procedural nature of these formulas allows statistical Taylor expansion to be implemented efficiently at the hardware level.
@@
A key open question is whether variance arithmetic can be adapted to achieve ideal coverage for floating-point rounding errors, because many theoretical calculations lack explicit input uncertainties. Variance arithmetic does not adjust uncertainty variance when floating-point rounding error occurs during calculation. This leads to error deviations larger than 1. Detecting floating-point rounding error also needs hardware implementation for efficiency.
@@
In variance arithmetic, deviations are comparable with values, but variances are used in calculation. This effectively limits the range of deviations to the square root of that of values. If the sign bit of the floating type can be re-purposed as an exponent bit in a new unsigned floating-point representation, the range of the deviation will be identical to that of values.
@@
When an analytic expression undergoes statistical Taylor expansion, the resulting expression can become highly complex, as in the case of matrix inversion. However, modern symbolic computation tools such as SymPy and Mathematica can greatly facilitate these calculations. This observation suggests that it may be time to shift from purely numerical programming toward analytic programming, particularly for problems that possess inherently analytic formulations.
@@
As an enhancement to dependency tracing, source tracing identifies each input’s contribution to the overall result uncertainty. This capability enables engineers to pinpoint the primary sources of measurement inaccuracy and in turn guide targeted improvements in data acquisition and processing strategies. For example, Formula (2.31) can guide how to improve the ideal ratio of $x\pm y$.
@@
Because traditional numerical approaches are based on floating-point arithmetic, they must be reexamined or even reinvented within the framework of variance arithmetic. For instance, most conventional numerical algorithms aim to identify optimal computational paths, whereas statistical Taylor expansion conceptually rejects all path-dependent calculations. Reconciling these two paradigms may present a significant and ongoing challenge.

### v18

@@ 9.3 Testing Signals
For Sin or Cos signals, forward and reverse DFT transforms differ in their data prospective of time domain versus frequency domain:
  • The forward transformation converts a time-domain sine or cosine signal into a frequency-domain spectrum in which most values are zero, causing its uncertainties to grow more rapidly during cancellation.
  • In contrast, the reverse transformation spreads the precise frequency-domain spectrum (where most values are zero) back into a time-domain sine or cosine signal, causing its uncertainties to grow more slowly.
@@ 9.4 Trigonometric Library Errors
Figure 23: Difference between Library and Quart $\sin(x)$ (y-axis) for $x=2\pi j/2^{L},j=0,1\dots 2^{L+2}$ (x-axis), and $L=5,6$ (legend). The uncertainty of the Quart $\sin(x)$ is $\sin(x)$ ULP, which shows a periodicity of $\pi$.
@@
Formulas (9.1) and (9.2) restrict the use of $\sin(x)$ and $\cos(x)$ to $x=2\pi j/2^{L}$, where $L$ is the FFT order. To minimize numerical errors in computing $\sin(x)$, the following indexed sine can be used in place of standard library sine functions:
@@
    The values of $\sin(\pi j/2^{L}),j\in[0,2^{L-2}]$ are the library sine directly, whereas the values of $\sin(\pi j/2^{L}),j\in[2^{L-2},2^{L-1}]$ are computed from the library $\cos(\pi(2^{L-1}-j)/2^{L})$.
@@
Because the Quart sine function strictly preserves the symmetry and periodicity of the sine function, it provides better numerical accuracy than the Library sine function.
@@
  • Figure 24 shows the value difference between the Quart and Library $\cos(x)/\sin(x)$ also increases roughly linearly with $|x|$, but is $10^{2}$ times larger than that observed for $\sin(x)$. Therefore, the linear spectrum in Formula (9.3) contains significantly larger numerical errors when computed using the Library sine functions.
@@ 9.5 Using Quart Sine for Sin/Cos Signals
Figure 25: Error distributions of normalized errors of Sin/Cos signals for forward, reverse and round-trip transformations (legend) using the Quart sine function. The FFT order is $18$.
With the FFT order as the specific dimension, the error deviations obtained using the Quart sine function for forward and reverse transformations resemble those shown in Figure 14, whereas those for round-trip transformation are nearly identical to those shown in Figure 15, independent of the signal frequency or whether a Sin or Cos signals is used. Therefore, the results for Sin and Cos signals across all frequencies are pooled together for statistical analysis, under the unified category Sin/Cos signals. When the FFT order $L$ is less than 8, the error deviations oscillate around $1$ because of an insufficient sample count of $2^{L}$. Even data for forward and reverse transformations are drastically different, variance arithmetic works effectively in all cases.
When $L=18$ and $\delta x=0$, Figure 25 shows that the error distributions of Sin/Cos signal resemble Normal distributions, with an additional Delta-like distribution at $\tilde{z}=0$. The error distribution of the reverse transformation is structured on top of the Normal distribution, suggesting that the reverse transformation is more sensitive to numerical errors in the sine function.
@@ 9.6 Using Library Sine for Sin/Cos Signals
Figure 26: Error distributions of normalized errors of Sin/Cos signal for forward, reverse and round-trip transformations (legend) computed using the Library sine function. The FFT order is $18$.
@@
Figure 27: FFT value error spectrum of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the forward transformation. The legend distinguishes between uncertainty and value error. The x-axis represents the frequency index, and the y-axis represents both uncertainty and value error.
@@
Figure 28: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the reverse transformation. The legend distinguishes between uncertainty and value error. The x-axis represents the time index, and the y-axis represents both uncertainty and value error.
@@
Figure 29: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the round-trip transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the frequency index, and the y-axis represents both uncertainty and value error.
@@
Figure 30: Error deviation (z-axis) of the FFT reverse transformation of $sin(f\frac{2\pi}{2^{L}}j)$ versus frequency $f$ (x-axis) and FFT order $L$ (y-axis).
@@
With FFT order as the specific dimension, the error deviations obtained using the Library sine for the forward and reverse transformations resemble those shown in Figure 14, whereas those for the round-trip transformation are nearly identical to those shown in Figure 15. In addition, the error deviations are larger than $1$ when $\delta x<10^{-15}$ for the forward transformation and $\delta x<10^{-14}$ for the reverse transformation.
@@
When $\delta x=0$, the error deviations for the reverse transformation increase with the FFT order, to $6.2$ at FFT order $18$. As shown in Figure 23, the Library sine function contains more numerical errors, as a result, the error distribution for the reverse transformation obtained using the Library sine function in Figure 26 is more structured and broader than that shown in Figure 25 obtained using the Quart sine function, whereas the error distributions for the forward transformations are more similar. This difference is consistent with a larger error deviation $6.2>1$ for the reverse transformation, compared with the comparable error deviation $1.1>1$ for the forward transformation when $\delta x=0$.
@@
When $\delta x=0$, using the Library sine function, for a sine wave with a frequency of $3$, Figures 27, 28, and 29 present the value errors for the forward, reverse, and round-trip transformations, respectively. In the reverse transformation, the value errors exhibit a clear trend of increasing with the time index. These large value errors appear systematic rather than random and visually resemble a resonant pattern. Similar increases are observed at other frequencies and FFT orders, as well as in computational results obtained using mathematical libraries such as SciPy. In contrast, such resonance is absent from the round-trip transformation shown in Figure 29, as well as when using the Quart sine function. Figure 30 demonstrates that the error deviations increase with sine or cosine frequency, regardless of FFT order $L$ when $L>8$. Figure 23 indicates that the numerical errors of $cos(x)/sin(x)$ obtained using the Library sine function increase with a periodicity of $\pi$, which can resonate with a signal whose periodicity is an integer multiply of $\pi$, producing the resonant pattern shown in Figure 28. At higher frequency, the resonant beats between the signal and the numerical errors in the Library sine function become stronger. To suppress this numerical error resonance, an input noise of $\delta x=10^{-14}$ must be added to the sine or cosine signals. Such resonance of numerical errors can easily and mistakenly be taken as signals.
9.7 Using Quart Sine for Linear Signals
@@
Figure 33: Error distributions of normalized errors of Linear signals for forward, reverse and round-trip transformations (legend) computed using the Quart sine function. The FFT order is $18$.
Figures 31 and 32 show the error deviations for the forward and the reverse transformations, respectively. The forward transformation exhibits a larger ideal coverage area than the reverse transformations: $\delta x>10^{-12}$ for the forward transformation, and $\delta x>10^{-8}$ for the reverse transformation. In other areas, both transformations achieve proper coverage with error values around $1$.
When $L=18$ and $\delta x=0$, the error distribution of the reverse transformation in Figure 33 is narrower than that shown in Figure 26. The corresponding error deviations are $1.5<6.2$, respectively.
The error deviations for the round-trip transformation resemble those in Figure 15 but with the FFT order as the specific dimension.
9.8 Using Library Sine for Linear Signals
Figure 34: Error distributions of normalized errors of Linear signals for forward, reverse and round-trip transformations (legend) computed using the Library sine function. The FFT order is $18$.
@@ 9.8 Using Library Sine for Linear Signals
Figure 37: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the round-trip transformations of Linear signals computed using the Library sine function.
Figure 34 shows that the reverse error distribution when $\delta x=0$ is no longer bound. The difference between Figures 34 and 33 is consistent with the large numerical errors as demonstrated in Figure 24. Variance arithmetic fails because of the large amount of unspecified numerical errors from the Library sine function.
Figures 35 and 36 show much smaller ideal coverage areas than those shown in Figures 31 and 32. Because uncertainty deviations grow more slowly in the reverse transformation than in the forward transformation, the former exhibits a smaller ideal coverage region. Outside of the ideal coverage region, proper coverage cannot be achieved for the reverse transformation. Furthermore, the range of input noise that produces ideal coverage decreases with increasing FFT order. At sufficiently high FFT orders (visually beyond FFT order $25$ for the reverse transformation), ideal coverage may no longer be achievable. Although FFT is widely regarded as one of the most robust numerical algorithms [5][10], and is generally insensitive to input errors, it can still fail because heavy calculation can amplify numerical errors in the Library sine function to completely obscure true signals. Such deterioration in calculation is not easily detectable when using conventional floating-point arithmetic.
Figure 37 shows that, even when variance arithmetic can no longer effectively track the value errors for either the forward or the reverse transformations, it can still effectively track the value errors for the round-trip transformation, as shown by the plateau region at high $L$ and low $\delta x$. Such error cancellation arises from dependence tracing in statistical Taylor expansion.
@@ 9.9 Ideal Coverage
Figure 38: Error deviation (left y-axis) and uncertainty deviation (right y-axis) of Linear signals versus FFT order (x-axis) and transformation types (legend) computed using the Library sine function.
Adding noise to the input can suppress unspecified input errors. After adding a Gaussian input noise of $\delta x=10^{-3}$ to a Linear signal when using the Library sine function, the error distributions for both the forward and the reverse transformations become Normal, whereas the error distribution for the round-trip transformation becomes Delta. Figure 38 illustrates the corresponding error deviations and uncertainty deviations versus FFT order:
@@
  • As expected, the resulting uncertainty deviations for the round-trip transformations remain equal to the corresponding input uncertainties of $10^{-3}$.
@@
  • As expected, the resulting error deviations for the round-trip transformations are far less than $1$ but increase exponentially with FFT order $L$ because of increasing calculation.
@@
Table 2 shows the minimal required noise to achieve ideal coverage for FFT transformations at FFT order $L=18$ for different signals and sine functions, which is consistent with the corresponding error distributions shown in Figures 25, 26, 33, and 34. An error distribution can reflect its input uncertainty coverage. Without knowing the precise result, similar histogram can be constructed from the result data set for $f$ using the calculated mean $\overline{f}$ and deviation $\delta f$. It is worth investigating if such an empirical histogram has similar power to reveal input uncertainty coverage.
10 Regressive Generation of Sine
Figure 39: The resulting precision (left y-axis) and uncertainties (right y-axis) of $\sin(\pi j/2^{18})$ versus index $j$ (x-axis) using either Quart or regressive sine function (legend).
$\sin(0)=\cos(\frac{\pi}{2})=0;\;\;\;\sin(\frac{\pi}{2})=\cos(0)=1;$ (10.1)
Formulas (10.2) and (10.3) calculate $\sin(\pi j/2^{L}),\cos(\pi j/2^{L}),j=0\dots 2^{L-2}$ repressively for regression order $L=0\dots 17$ starting from Formula (10.1). Formula (10.4) shows that such regression guarantees both $\sin(x)^{2}+\cos(x)^{2}=1$ and $\sin(2x)=2\sin(x)\cos(x)$, such that value errors will not accumulate when the regression order $L$ increases.
Formula (10.2) is not suitable for computing $\sin(x)$ as $x\rightarrow 0$ because it suffers from behavior analogous to catastrophic cancellation [9][10]. As shown in Figure 39, the Quart sine function exhibits a constant precision around $10^{-16}$, whereas the regression sine function shows increasing precision up to $10^{-7}$ as $x\rightarrow 0$. Unlike the sneaky catastrophic cancellation in floating-point arithmetic, variance arithmetic uses coarser precision to demonstrate where and by how much the regression algorithm becomes unfit to compute $\sin(x)$.
@@ 11.1 Summary
When the uncorrelated uncertainty condition is satisfied, statistical Taylor expansion produces the mean, deviation, and reliability of an analytic expression. It tracks the variable dependencies in intermediate steps and rejects invalid calculations. Unlike conventional approaches, it explicitly incorporates the sample counts and uncertainty distributions into its result. Although statistical Taylor expansion eliminates the dependency problem, it also reduces execution flexibility.
The presence of ideal coverage is a necessary condition for a numerical algorithm based on statistical Taylor expansion to be considered correct. Ideal coverage defines the optimal range of applicability for an algorithm.
Variance arithmetic simplifies statistical Taylor expansion by introducing numerical rules that eliminate invalid results, such as divergent, negative variance, unstable, infinite, or unreliable results. It also provides proper coverage for floating-point rounding errors. The applicability of variance arithmetic has been demonstrated across a wide range of computational scenarios.
The code and analysis framework for variance arithmetic are available as an open-source project at https://github.com/Chengpu0707/VarianceArithmetic. A more detailed description of this study is presented at https://arxiv.org/abs/2410.01223.
@@ 11.2 Improvements Needed
This study presents statistical Taylor expansion and variance arithmetic, which are still in early stages of development. Accordingly, several important questions remain.
Library mathematical functions should be recalculated using variance arithmetic, to ensure that each output value is accompanied by its corresponding uncertainty. Without this refinement, the value errors in the library functions can produce unpredictable and potentially significant result errors.
@@
Bound momentum $\zeta(n,\kappa)$ should be extended to all probability distributions. The choice of ideal bounding range $\hat{\kappa}$ should be extended to other distributions, perhaps by using the same value of $1-\zeta(2,\kappa)$. The procedure to find the bound moments $\kappa$ for each sample count $N$ should be developed for discrete distributions. The statistical meaning for ideal leakage $\varepsilon\in[0,1]$ in the context of its choice of ideal bounding range $\hat{\kappa}$ needs further investigation: Is $\varepsilon$ the possibility for the result to be not bound by $\hat{\kappa}$?
@@
The performance of variance arithmetic must be improved for broader practical adoption. The fundamental formulas of statistical Taylor expansion, Formulas (2.5), (2.3), (2.8), and (2.3), contain many independent summations, making them excellent candidates for parallel processing. Moreover, the inherently procedural nature of these formulas allows statistical Taylor expansion to be implemented efficiently at the hardware level.
@@
A key open question is whether variance arithmetic can be adapted to achieve ideal coverage for floating-point rounding errors, because many theoretical calculations lack explicit input uncertainties. Variance arithmetic does not adjust uncertainty characterization when floating-point rounding error occurs during calculation, which leads to error deviations larger than 1. Detecting floating-point rounding error and adjusting uncertainty characterization in real time needs hardware implementation for efficiency.
@@
In variance arithmetic, deviations are comparable to values; however variances are used in calculation. This approach effectively limits the range of deviations to the square root of that of the values. If the sign bit of the floating type can be re-purposed as an exponent bit in a new unsigned floating-point representation, the range of the deviations will be identical to that of the values.
@@
When an analytic expression undergoes statistical Taylor expansion, the resulting expression can become highly complex, as in the case of matrix inversion. Modern symbolic computation tools such as SymPy and Mathematica can significantly facilitate such calculations. This observation suggests that it may be time to shift from purely numerical programming toward analytic programming, particularly for problems that possess inherently analytic formulations.
@@
As an enhancement to dependency tracing, source tracing identifies each input’s contribution to the overall result uncertainty. This capability enables engineers to pinpoint the primary sources of measurement inaccuracy and guides them to targeted improvements in data acquisition and processing strategies. For example, Formula (2.33) can provide guidance on improving the ideal leakage of $x\pm y$.
@@
Because conventional numerical approaches are based on floating-point arithmetic and path dependent in general, they must be reexamined or even reinvented within the framework of variance arithmetic. For instance, most conventional numerical algorithms aim to identify optimal computational paths, whereas statistical Taylor expansion conceptually rejects all path-dependent calculations. Reconciling these two paradigms may present a significant and ongoing challenge.